OpenStudy (atomicreaper):

What are the roots of 3x2 + 10 = 4x?

OpenStudy (jango_in_dtown):

3x^2+10=4x impies 3x^2-4x+10=0 by applying Sridhar Acharya's method roots re \[(2/3)\pm (\sqrt{26}/3)i\]